TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x06x21x19x42

Powering Tests Sector 7-8 - Report 21.06.2007

Test Report of Thursday 21.06.2007

  • RQD.A78
    • PLI3.2 Discharge Request - Test OK
    • PLI3.3 Quench heater discharge in 12R7 - Test OK - Cryo_Maintain was recovered after 10 minutes and Cryo_Start in about 30 minutes.

  • RQF.A78
    • PIC2 completed - Test OK
    • PLI1.1 - Current cycle at 760A completed - Test OK
    • PLI1.3 - PC Failure at 760A completed - Test OK
    • Communication problems with the energy extraction controller prevented from closing the switches remotely. Tests stopped.

  • RB.A78
    • QPS IST completed - 2 DQHDS had to be replaced
    • A communication problem came up due to a corrupted code line maybe caused during the downloading of the last software version.
    • Current leads connected to DC cables in DFBAO and DFBAN
    • Last corrections being done by QPS experts

  • Matching Section
    • Heat runs continue during the whole week with RD2 and RQ5
    • Type tests of the converters by AB/PO completed.

Planning for Friday 22.06.2007

  • RB.A78
    • PIC1
    • Connection of the 60A conveter and TF measurement
    • Connection of the RB converter to the circuit+ PCC
    • PIC2

  • RQF.A78 (in parallel with RB whenever possible)
    • Complete PLI1 (760A)
      • Energy extraction discharge
      • FPA
      • Discharge request from PC
      • Heater firing
    • PLI2 (2000A)
    • PLI3 (6500A)

  • 60 Closed Orbit Correctors
    • PCC and P2N on selected circuits

  • AB/PO current stability measurements in RD2.L8

  • Empty of DFBMA and Q4/D2 cryostats over the weekend

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-06-22 - AntonioVergara
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login